### v2

# DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding

Abstract
Text-rich document understanding (TDU) refers to analyzing and comprehending documents containing substantial textual content. With the rapid evolution of large language models (LLMs), they have been widely leveraged for TDU due to their remarkable versatility and generalization. In this paper, we introduce DocLayLLM, an efficient and effective multi-modal extension of LLMs specifically designed for TDU. By integrating visual patch tokens and 2D positional tokens into LLMs and encoding the document content using the LLMs themselves, we fully take advantage of the document comprehension capability of LLMs and enhance their perception of OCR information. We have also deeply considered the role of the chain-of-thought (CoT) and innovatively proposed the techniques of CoT Pre-training and CoT Annealing. Our DocLayLLM can achieve remarkable performances with lightweight training settings, showcasing its efficiency and effectiveness. Experimental results demonstrate that our DocLayLLM surpasses existing OCR-dependent methods and also outperforms OCR-free competitors.
DocLayLLM: An Efficient and Effective Multi-modal Extension of Large Language Models for Text-rich Document Understanding
Wenhui Liao1, Jiapeng Wang1, Hongliang Li1, Chengyu Wang2, Jun Huang2, Lianwen Jin1 1South China University of Technology, Guangzhou, China 2Alibaba Group, Hangzhou, China {eelwh, eejpwang, eehongliangli}@mail.scut.edu.cn, eelwjin@scut.edu.cn {chengyu.wcy, huangjun.hj}@alibaba-inc.com
22footnotetext: Equal contributions, listed in no particular order.11footnotetext: Co-corresponding authors.
1 Introduction
Text-rich document understanding (TDU) involves analyzing and comprehending documents with extensive textual information. These documents often have complex layouts, requiring an integrated understanding of text, layout, and image data. Early research (Xu et al., 2020; Li et al., 2021a; Appalaraju et al., 2021; Li et al., 2021b; Wang et al., 2022; Hong et al., 2022; Luo et al., 2023) often used BERT-like (Devlin et al., 2019) Transformer encoders, designing various pre-training tasks for large-scale unlabeled document corpora and fine-tuning their models on specific downstream tasks like key information extraction (KIE) and visual question answering (VQA). However, these pre-training tasks often diverge significantly from downstream tasks, requiring the model to acquire substantial task-specific knowledge during fine-tuning. This may result in limited generalization capability across different tasks and datasets.
With the rapid development of large language models (LLMs), their impressive versatility and generalization capabilities have captured widespread interest. This has led many researchers to explore LLMs to overcome the challenges of TDU (Zhang et al., 2023; Ye et al., 2023b; Tang et al., 2024; Hu et al., 2024a; Li et al., 2024b). Initially, efforts focused on employing multi-modal large language models (MLLMs) for OCR-free TDU, where document images are directly fed into MLLMs, generating responses to comprehension tasks. These MLLMs, which can process images and text inputs effectively, are among the most recent advancements in this domain. To date, several general MLLMs have been developed to manage diverse scenarios. Researchers often enhance these models by integrating additional document instruction-tuning data, aiming to improve their ability to comprehend documents. However, these models typically necessitate high-resolution images and an amount of training data to recognize text accurately in dense documents, which escalates the computational demands. Furthermore, to achieve superior performance, these models are usually trained using vast collections of image and instruction data, including the training sets of a specific dataset they are tested on, potentially affecting the evaluation of their ability to perform on previously unseen datasets.
To overcome the constraints of OCR-free methods, scholars have pursued diverse methodologies for integrating OCR information into LLMs. Initial strategies (Wang et al., 2023; He et al., 2023) involved inputting coordinates of OCR bounding boxes as numeric text. However, this textual representation of numbers often resulted in excessively long inputs that could decrease inference speeds and potentially degrade model performance. In response, many researchers Kim et al. (2023); Fujitake (2024); Luo et al. (2024) have adopted an auxiliary document encoder (DE) like LayoutLMv3 (Huang et al., 2022) to process textual, visual, and layout features from documents and integrated into LLMs. Although these methods avoid high-resolution image inputs, the textual content of OCR documents must also be encoded alongside the layout and image features with the DE due to its structural limitations. However, the document comprehension capabilities of these DEs may not be on par with those of LLMs themselves, which might have led to significant fine-tuning efforts for aligning the feature spaces of DEs and LLMs during the whole training phase, demanding substantial computational resources.
In response to the limitations inherent in OCR-dependent TDU, we introduce DocLayLLM, an efficient and effective multimodal extension of large language models for text-rich document understanding. Our method diverges from existing approaches that utilize an extra DE by integrating image patch tokens and 2D positional tokens with textual content into a natural language expression and processing them through an LLM. This innovative integration not only taps into the LLM’s robust text comprehension abilities but also substantially enhances its capability to perceive document layouts. Besides, by incorporating patch tokens, the model gains a rudimentary understanding of document structures, and supports image-related pre-training tasks.
To advance model performance, we have fully integrated the concept of the chain-of-thought (CoT) (Brown et al., 2020) into every stage of our training process. Firstly, we developed a novel CoT Pre-training strategy. CoT is engineered to preserve thematic coherence and logical progression during model reasoning, further enhancing the inference ability of LLM. It has demonstrated effectiveness in various subsequent studies (Wei et al., 2022; He et al., 2023; Luo et al., 2024). Nevertheless, the incorporation of CoT in the pre-training phase to boost reasoning capabilities remains underexplored. By employing CoT, we have improved the model’s comprehension ability while aligning these visual and layout features with LLM features.
Furthermore, we introduced a CoT Annealing technique. While previous research (Li et al., 2024a; Llama Team, 2024) has validated the effectiveness of annealing strategies that gradually increase the proportion of high-quality data towards the end of training, the impact of CoT on data quality had not been explored. To address this, we first propose an annealing strategy that considers data quality from a CoT perspective, further improving overall model performance.
Thanks to the above design, utilizing only around 3M pre-training data and 300K supervised fine-tuning (SFT) data with an efficient tuning technology of LoRA (Hu et al., 2022), our DocLayLLM finishes the whole training process in 36 A100 days, while some TDU model could take up to 1200 A100 days (Tang et al., 2024), therefore resulting a highly lightweight extension of LLM. However, even with such low resources, our DocLayLLM surpasses existing OCR-dependent TDU methods. Furthermore, it also excels in specific tasks without fine-tuning on downstream tasks and outperforms state-of-the-art (SOTA) OCR-free methods under fair fine-tuning conditions, showcasing our efficiency and effectiveness. Our contributions are:
  • We propose an efficient and effective multi-modal extension of LLMs, which has augmented LLM in understanding text-rich documents and significantly reduced training resources in need.
  • We introduce CoT Pre-training, remarkably enhancing the model’s data utilization efficiency and comprehension abilities.
  • We re-evaluate data annealing from a novel perspective and introduce the CoT Annealing strategy, which further enhances the model’s efficiency in data utilization and boosts its ultimate performance.
  • Our method outperforms existing OCR-dependent TDU solutions and surpasses current OCR-free methods under comparable supervised fine-tuning conditions.
2 Related Works
OCR-free MLLMs for Text-Rich Documents Understanding.  Many studies have attempted to directly answer the instruction for TDU tasks from document images, thereby obviating the need for OCR processing and resulting in an end-to-end paradigm. Such models are referred to as OCR-free models. Some early examples like Donut (Kim et al., 2022), UDOP (Tang et al., 2023), and OmniParser (Wan et al., 2024) have explored end-to-end document understanding solutions without relying on LLMs or MLLMs. With the rapid development of MLLMs, many new studies have begun to focus on enhancing the capabilities of general MLLMs in document understanding. Among these, LLaVAR (Zhang et al., 2023) is an early initiative that studied OCR-free MLLMs, utilizing GPT-4 (OpenAI, 2023) to generate image-text dialogue data for fine-tuning the LLaVA (Liu et al., 2023a) model. Inspired by LLaVAR, TextSquare (Tang et al., 2024) further uses Gemini (Google, 2023) to create 10M visual dialogue data covering more text-rich and other scenarios to augmente InternLM-Xcomposer2 (Dong et al., 2024). In addition, mPLUG-DocOwl 1.5 (Hu et al., 2024a) enhances the document understanding capability by integrating existing document comprehension data and converting it into vision-related dialogue data. Moreover, researchers have also explored in terms of image encoding and image-text alignment in MLLMs. For instance, UReader Ye et al. (2023b) and Monkey Liu et al. (2024) empower MLLMs to handle high-resolution images by segmenting these images into multiple sub-images. Then they concatenate the independently encoded features of these sub-images along with a thumbnail of the original image before feeding them into the LLM. Although these methods have achieved notable success in document understanding tasks, they typically require high-resolution image inputs to identify text within images, significantly increasing computational demands. Additionally, the absence of explicit OCR information input means that these methods still have room for improvement in tasks highly dependent on OCRed positional information like KIE. Besides, to further enhance performance, these methods often train using extensive datasets, including the training sets of the datasets used for evaluation.
Figure 1: The overall architecture of DocLayLLM.
OCR-dependent MLLMs for Text-Rich Documents Understanding.  To address the challenges of OCR-free MLLMs, some researchers have attempted to integrate OCR information from text-rich documents into LLMs to enhance their perception of OCR information. ICL-D3IE (He et al., 2023) feeds coordinates of OCR boxes into the LLM using the format of [left, top, right, bottom], while LATIN-Prompt (Wang et al., 2023) aims to model the document layout by inserting spaces and newline characters between lines of OCR-recognized text, thus providing the LLM with a document input containing rudimentary layout information. However, these methods of inputting OCR information lead to overly lengthy inputs, subsequently slowing inference and impacting model performance. Consequently, subsequent studies have explored encoding OCR information using additional encoders. Cream (Kim et al., 2023) utilizes BLIP-2 (Li et al., 2023) to encode document OCR positions, images, text, and additional object detection information, with the output of BLIP-2 fed directly into the LLM. InstructDoc (Tanaka et al., 2024) introduces an extra feature extractor to encode document OCR information, compressing it with learnable tokens before inputting it into the LLM. Additionally, both (Fujitake, 2024) and (Luo et al., 2024) adopt LayoutLMv3 (Huang et al., 2022) to encode the OCR information of documents before inputting them into the LLM. Although these methods avoid the issue of overly lengthy text inputs, they must use the extra DE to encode all the OCR textual, positional information of the document due to the structural restriction of DE. However, the additional DE might not be comparable with LLMs themselves in terms of the document comprehension capability. This might be the reason why they usually need extensive training of both the DE and the LLM to align the features encoded by the DE with those of the LLM, which consumes significant computational resources. DocLLM (Wang et al., 2024) proposes an extension where the hidden states derived from encoded OCR positional information interact with the hidden states of OCR text features encoded by LLMs within each Transformer block via cross-attention. However, as it alters the internal structure of the LLM, it could affect its generality and generalizability, thus necessitating fine-tuning before deployment. Moreover, it does not adequately consider the role of image features in TDU. Overall, designing an efficient and effective method to provide OCR information to LLMs remains a research question worthy of deep exploration.
3 DocLayLLM
DocLayLLM is an LLM-based methodology that encodes OCRed textual, visual, and positional information directly within an LLM, eliminating the need for additional DE. The training regime of DocLayLLM includes both pre-training and supervised fine-tuning phases, which are both instruction-tuning-based processes. To further enhance our model, we propose the integration of the CoT into our pre-training tasks, denoted as CoT Pre-training. Additionally, we introduce CoT Annealing, which involves the application of annealing to the data utilized in the supervised fine-tuning phase, specifically from the perspective of CoT.
3.1 Model Architecture
The overall architecture of our DocLayLLM is illustrated in Figure 1. To process a document image, we initially employ an OCR engine. This engine extracts the OCR results, which encompass the textual content, denoted as $Text_{0:N}$, and the associated bounding box, represented by $Box_{0:N}$. Here, $N$ indicates the total number of text lines detected by the OCR, and each $Box_{i}=[x_{0},y_{0},x_{1},y_{1}]$ defines the bounding box for $Text_{i}$. The bounding box coordinates correspond to the left, top, right, and bottom boundaries of the text line. After the OCR process, the document image is segmented into patches, denoted as $Patch_{0:M}$, where $M$ represents the total number of patches. Next, the OCR results and the document patches are integrated with the fixed part of the prompt and the specific $Question$ related to the document. This amalgamation forms the final prompt input to the model, presented in a natural language expression $\mathcal{T}$: Given the document patches: {$Patch_{0:M}$} and the document text contents and locations in the form of “text, [left, top, right, bottom]: {$(Text_{i},Box_{i})_{0:N}$} {$Question$}
Subsequently, the textual components of the prompt, including the fixed expression, $Text_{0:N}$, and $Question$, are processed using the LLM’s native text embedder. This generates the embeddings for the text content as follows:
$Emb_{T}=TE(TextContent),$ (1)
where $TE$ is the native text embedder, and $TextContent$ consists of the fixed expression, $Text_{0:N}$, and $Question$.
As for the $Patch_{0:M}$ and $Box_{0:N}$, they are first embedded with an extra embedder, then projected to the LLM’s feature space via a linear projector to obtain $Emb_{V}$ and $Emb_{L}$ as:
$Emb_{V}$ $=VP(VE(Patch_{0:M})),$ (2)
$Emb_{L}$ $=LP(LE(Box_{0:N})),$ (3)
where $VE$ and $VP$ represent the visual embedder and visual projector, $LE$ and $LP$ denote the layout embedder and layout projector, respectively. It is worth noting that despite the introduction of $VE$, $VP$, $LE$, and $LP$, these are all simple embedding layers or linear layers, thereby minimizing additional computational burden.
Finally, LLM will comprehend the embeddings provided and utilize them to generate the $Answer$ to the $Question$:
$\begin{split}Answer=LLM(\pi((Emb_{T},Emb_{V},Emb_{L}),\mathcal{T})).\end{split}$ (4)
Here $\pi$ represents organizing $Emb_{T}$, $Emb_{V}$, $Emb_{L}$ into a feature sequence according to $\mathcal{T}$.
Through our design, we have preserved the original structure of the LLM, ensuring its generalization capability. At the same time, we have enabled the LLM to easily accept and process visual and layout features, making it better suited for understanding text-rich documents with various layout. This approach results in a simple yet efficient and effective multi-modal extension.
Task Description
Document Description Provide a brief overview of the document.
Text and Box Reconstruction Recover the coordinates of bounding boxes of all the OCR text.
Layout Analysis Determine the layout type (e.g., Title, Figure, Paragraph, etc) and locate specific layout elements.
Table Analysis Decode the structure of tables and identify the positions of elements within.
Mask Language Model Restore masked words in the provided OCR text.
Mask Position Model Reconstruct the box for text elements missing the coordinates of the bounding box.
Geometric Analysis Calculate distances or directions between two specified text elements.
Table 1: The training tasks during the pre-training stage.
3.2 CoT Pre-training
During the pre-training phase, we incorporate a wide range of tasks (as detailed in Table 1) to enhance the LLM’s understanding of the visual and layout information inspired by (Luo et al., 2024). Since the datasets related to these tasks are not designed explicitly for instruction-tuning, their annotations are not in a QA format. Therefore, we manually create instructions based on these annotations for training. However, simply converting task descriptions into $Question$ and annotations into $Answer$ is highly inefficient regarding data utilization. To address this, we incorporate a CoT approach during the conversion to instruction-tuning data for specific tasks, asking LLM to solve our problem step-by-step, enhancing the model’s comprehension abilities. Our designed CoT primarily falls into the following three categories:
3.2.1 Layout-Type-Related CoT
We conduct a layout-type-related CoT for the layout analysis task. To determine the layout type of a given area, it is beneficial first to locate it within the document and then infer based on the nearby text lines and the overall document layout. When giving an area bounded by a bounding box, our layout-type-related CoT can be formulated as following steps:
  1. 1.
    Identify the text within the area and, based on this, determine its approximate location within the document (e.g., top left, top right part of the document, etc).
  2. 2.
    Locate the nearest bounding boxes and assess their layout type.
  3. 3.
    Infer the layout type of the given area based on the above information.
Through this CoT process, we enable the LLM to fully utilize the visual, layout, and textual features of the document, thereby improving their ability to understand the document’s overall structure.
3.2.2 Table-Structure-Aware CoT
Table-structure-aware CoT is primarily employed in the table analysis task. We have integrated the XY-Cut algorithm Nagy et al. (1992)—a widely-used technique for analyzing table structures. For instance, to locate an element at the $i$-th row and $j$-th column, the CoT process is outlined as follows:
  1. 1.
    Identify all the table heads and their respective locations. The head of the $j$-th column corresponds to the $j$-th element in the heads list.
  2. 2.
    Output all elements in the $j$-th column when told that the elements of the same column are generally vertically aligned with their corresponding heads.
  3. 3.
    Determine the final answer based on the $i$-th element of the contents in the $j$-th column.
Applying the above CoT, we break down a complex question into simple steps using the basic XY-Cut thought. This approach improves the LLM’s inference ability through this breakdown and enhances data utilization by effectively leveraging the original annotations.
3.2.3 Geometry-Related CoT
This is used in geometric analysis, targeting the identification of geometric relationships between two text lines. It improves LLM’s understanding of layout features and enhances inference with geometric relations. Specifically, when giving the text content of two lines:
  1. 1.
    Recover the bounding box of the given text lines.
  2. 2.
    Analyze their vertical and horizontal projections to determine if they overlap through interval analysis.
  3. 3.
    Calculate the center point of each text line and determine their relative location by comparing their center points.
  4. 4.
    Calculate the minimum Euclidean distance based on the previous analysis. If they overlap, the distance is 0; if they overlap in vertical/horizontal projections, the minimum distance is the smallest interval in the horizontal/vertical projections; otherwise, it is the minimum Euclidean distance between their nearest corners.
This design enhances the LLM’s arithmetic reasoning capabilities and helps it more accurately grasp the spatial relationships between different boxes, and also contributes to layout-type-aware CoT.
3.3 CoT Annealing
Data annealing involves gradually increasing the proportion of high-quality data towards the end of supervised fine-tuning to improve data utilization efficiency and enhance the model’s reasoning abilities. This technique has been proven highly effective and is widely adopted in models like MiniCPM Hu et al. (2024b), Data-CompLM (Li et al., 2024a), and Llama-3.1 (Llama Team, 2024).
During the SFT stage, we employ data with a layout-aware CoT Luo et al. (2024) that helps focus on potential regions of questions and answers within documents, thereby better leveraging the layout features learned during pre-training. It also allows the model to enhance its reasoning ability through step-by-step thinking. However, in the latter stages of SFT, our goal for LLM is to output precise and direct answers to questions. At this point, using CoT data may introduce unnecessary noise with a long response, and the overly uniform layout-aware CoT data might partially compromise the LLM’s generalization abilities.
To balance the strengths and weaknesses of such data, we propose CoT annealing, which reconsiders this data annealing issue from a CoT perspective. Specifically, we generate direct-answer data without CoT from the CoT data, treating it as high-quality data. During the SFT process, we gradually increase the proportion of this direct-answer data, thereby enhancing the LLM’s ability to respond with a clear and direct answer. We initially utilize only the data with CoT, gradually adjusting the ratio of CoT to non-CoT data as the SFT process progresses. Ultimately, we conclude the training using only the data without CoT.
Method Document-oriented VQA KIE
DocVQA VisualMRC Avg. FUNSD- CORD- SROIE- Avg.
Text
Llama2-7B-Chat (Touvron et al., 2023) 64.99 52.84 58.92 48.20 47.70 68.97 54.96
Llama3-8B-Instruct (Llama Team, 2024) 51.79 47.77 49.78 68.57 52.31 61.24 60.70
Text + Layout
Llama2-7B-Chatlayout (Touvron et al., 2023) 56.55 49.26 52.91 58.34 50.93 51.15 53.47
Llama3-8B-Instructlayout (Llama Team, 2024) 49.13 41.75 45.44 74.00 62.20 63.15 66.45
LayTextLLMzero (Lu et al., 2024) 65.50 37.40 51.45 72.00 45.50 82.00 66.50
Text + Layout + Image
LayoutLLMLlama2 (Luo et al., 2024) 74.25 55.73 64.99 78.65 62.21 70.97 70.61
DocLayLLMLlama2 (Ours) 72.83 55.04 63.94 78.67 70.81 83.27 77.58
DocLayLLMLlama3 (Ours) 78.40 58.55 68.48 84.12 71.34 84.36 79.94
Table 2: Comparison with OCR-dependent methods using metrics from LayoutLLM Luo et al. (2024). Avg. represents the average performance across different benchmarks, and - indicates the cleaned test set used in LayoutLLM. The suffix layout denotes the method from He et al. (2023), which provides LLM with layout features via OCR box coordinates in textual form. Bold indicates the best performance, and underline indicates the second-best one.
Method Document-oriented VQA KIE
DocVQA VisualMRC Avg. FUNSD CORD SROIE Avg.
Text
Llama2-7B-Chat (Touvron et al., 2023) 20.50 6.30 13.40 23.40 51.80 58.60 44.60
Llama3-8B-Instruct (Llama Team, 2024) 51.79 229.74 140.77 66.67 74.71 82.51 74.63
Text + Layout
Llama2-7B-Chatlayout (Touvron et al., 2023) 12.30 28.00 20.15 14.40 38.10 50.60 34.37
Llama3-8B-Instructlayout (Llama Team, 2024) 49.13 211.69 130.41 74.71 75.33 75.93 75.32
DocLLM (Wang et al., 2024) 69.50 264.10 166.80 51.80 67.60 91.90 70.43
LayTextLLMzero (Lu et al., 2024) 65.50 200.20 132.85 47.20 77.20 83.80 69.40
Text + Layout + Image
DocLayLLMLlama2 (Ours) 72.83 310.60 191.72 80.74 79.37 84.37 81.49
DocLayLLMLlama3 (Ours) 78.40 357.04 217.72 86.47 83.66 86.08 85.40
Table 3: Comparison with OCR-dependent methods using metrics from DocLLM Wang et al. (2024). indicates that the model uses the training set corresponding to the evaluation set during training.
4 Experiments
4.1 Datasets
During the pre-training stage, we conduct our pre-training tasks on several datasets, including RVL-CDIP (Pramanik et al., 2020), PubLayNet (Zhong et al., 2019), PubTabNet (Zhong et al., 2020), DocLayNet (Pfitzmann et al., 2022), DocBank (Li et al., 2020), DocILE (Simsa et al., 2023), and the Document Dense Description dataset from LayoutLLM (Luo et al., 2024). This results in a total of 3.1 million pairs of in-context learning data, which is significantly less than the volume used in previous OCR-dependent TDU methods. For the supervised fine-tuning stage, we apply CoT annealing on the layout-aware document QA data utilized in LayoutLLM (Luo et al., 2024).
We evaluate our DocLayLLM on widely used text-rich document benchmarks, including FUNSD (Jaume et al., 2019), CORD (Park et al., 2019), SROIE (Huang et al., 2019), POIE (Kuang et al., 2023), DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), VisualMRC (Tanaka et al., 2021), DeepForm (Svetlichnaya, 2020), KLC (Stanislawek et al., 2021), WTQ (Pasupat and Liang, 2015), and TabFact (Chen et al., 2020). To fairly compare DocLayLLM with other approaches, we follow the evaluation pipeline and metrics they used.
4.2 Implement Details
Our method is based on Llama3-8B-Instruct  [^1] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct (Llama Team, 2024). We also build a version using Llama2-7B-Chat  [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023) for comparison with other methods, denoted as DocLayLLMLlama3 and DocLayLLMLlama2 respectively. We initialize our visual encoder and layout encoder weights from LayoutLMv3 (Huang et al., 2022). During pre-training and fine-tuning, we use LoRA Hu et al. (2022) with a rank of 64. We train our model on 8 NVIDIA Tesla A100-80G GPUs, taking about 30 A100 days for pre-training and 6 A100 days for fine-tuning. For OCR results, we use the default OCR data from the official datasets or commercial OCR engines if not provided. Additional details can be found in Appendix A.1.
Method Document-oriented VQA KIE
DocVQA InfoVQA Avg. FUNSD SROIE POIE Avg.
LLaVAR (Zhang et al., 2023) 12.30 16.50 14.40 0.50 5.20 5.90 3.87
UniDoc (Feng et al., 2023b) 7.70 14.70 11.20 1.00 2.90 5.10 3.00
DocPedia (Feng et al., 2023a) 47.10 15.20 31.15 29.90 21.40 39.90 30.40
Monkey (Li et al., 2024b) 50.10 25.80 37.95 24.10 41.90 19.90 28.63
TextMonkey (Liu et al., 2024) 66.70 28.60 47.65 42.90 46.20 32.00 40.37
DocOwl 1.5 (Hu et al., 2024a) - - - - 48.30 51.80 -
TextSquare (Tang et al., 2024) - - - - 53.20 71.80 -
DocKylin (Zhang et al., 2024) 65.10 34.80 49.95 58.70 25.50 49.50 44.57
DocLayLLMLlama3 (Ours) 77.79 42.02 59.91 80.28 76.59 75.13 77.33
Table 4: Comparison with OCR-free methods using the precision metric from TextMonkey Liu et al. (2024). indicates that the model uses the training set corresponding to the evaluation set during training.
Method Document-oriented VQA KIE Table Understanding
DocVQA InfoVQA VisualMRC Avg. DeepForm KLC Avg. WTQ TabFact Avg.
DocPedia (Feng et al., 2023a) 47.10 15.20 - - - - - - - -
DocOwl (Ye et al., 2023a) 62.20 38.20 188.80 96.40 42.60 30.30 36.45 26.90 60.20 43.55
UReader (Ye et al., 2023b) 65.40 42.20 221.70 109.77 49.50 32.80 41.15 29.40 67.60 48.50
Monkey (Li et al., 2024b) 66.50 36.10 - - 40.60 32.80 - 25.30 - -
TextMonkey (Liu et al., 2024) 73.00 - - - 59.70 37.80 - 31.90 - -
DocKylin (Zhang et al., 2024) 77.30 46.60 319.90 147.93 64.20 - - 32.40 - -
CogAgent (Hong et al., 2024) 81.60 44.50 - - - - - - - -
DocOwl 1.5 (Hu et al., 2024a) 82.20 50.70 246.40 126.43 68.80 38.70 53.75 40.60 80.20 60.40
TextSquare (Tang et al., 2024) 84.30 51.50 - - - - 49.70 84.20 66.95
DocLayLLMLlama3 (Ours) 78.40 40.90 357.04 158.80 30.19 25.76 27.98 46.03 65.65 55.84
DocLayLLMLlama3 (Ours) 86.52 58.36 388.85 177.91 77.07 40.73 58.90 58.62 83.36 70.99
Table 5: Comparison with OCR-free methods using metrics from DocOwl 1.5 Ye et al. (2023a). indicates that the model uses the training set corresponding to the evaluation set during training.
4.3 Comparisons with SOTA OCR-dependent methods
We have compared our DocLayLLM with existing SOTA OCR-dependent methods. In order to exhibit the zero-shot inferring ability of OCR-dependent methods precisely, we chose their performances under the zero-shot scenario unless they are not provided. The experimental results are shown in Table 2 and Table 3.
It can be found that our method outperforms the baseline models, regardless of the type of baseline model used or whether the document’s layout information is provided to the baseline model. It validates the rationale behind our design of multi-modal extension.
Additionally, our method significantly outperforms previous SOTA methods, even with less pre-training data and fewer resources. As shown in Table 2, it surpasses the performance of LayoutLLM Luo et al. (2024) and LayTextLLM Lu et al. (2024) while using fewer computational resources by solely employing LoRA during training. Even with Llama2 as the backbone, our approach matches the performance of previous SOTA methods in the Document-oriented VQA task and significantly outperforms them in the KIE tasks, further demonstrating the efficiency and effectiveness of our approach.
Besides, our method exhibits robust zero-shot generalization capability without the need for fine-tuning on specific downstream tasks. From the comparison with DocLLM Wang et al. (2024) in Table 3, it is clear that our method, even without fine-tuning, significantly outperforms DocLLM across most datasets and achieves a considerable overall performance improvement. This validates our design, which preserves the strong generalization ability of the original LLM structure.
In summary, our DocLayLLM is an efficient and effective OCR-dependent TDU model, demonstrating powerful zero-shot generalization capability by encoding the OCR information of the document with LLM and setting the new SOTA performance compared to other OCR-dependent competitors.
4.4 Comparisons with SOTA OCR-free methods
We have also compared our method with SOTA OCR-free TDU models. Table 4 presents the precision of our method and some advent OCR-free TDU methods across different benchmarks. As shown in the table, even though in the circumstance that these OCR-free models typically use the corresponding training sets of these benchmarks during training, while our method infers in a purely zero-shot manner, our DocLayLLM still outperforms these models, particularly in the KIE task, which heavily relies on layout information. This clearly demonstrates the superiority of our design which provides LLM with OCR information.
We additionally test DocLayLLM on a broader set of benchmarks from DocOwl 1.5 Ye et al. (2023a). To further assess the performance of DocLayLLM on seen datasets, we incorporate the training sets of these downstream benchmarks used in DocOwl 1.5 during the SFT process. It is worth noting that OCR-free TDU methods typically include these datasets during training. To maintain the efficiency advantage of our method, we ensure that the SFT is conducted with the same number of steps as the standard DocLayLLM. The results in Table 5 show that with the inclusion of downstream task data during SFT, our method outperforms the SOTA OCR-free TDU models in most datasets. Moreover, our overall performance in the Document-oriented VQA, KIE, and Table Understanding sub-tasks surpasses the OCR-free TDU methods using limited training resources, showcasing the tremendous potential of OCR-dependent methods and further reinforcing the reliability of our method.
Overall, even with fewer training resources, our DocLayLLM shows strong zero-shot inference performance, outperforming many OCR-free TDU methods without prior exposure to downstream datasets. Additionally, when incorporating downstream datasets during SFT like other OCR-free approaches, DocLayLLM surpasses them, highlighting its potential in solving TDU tasks by integrating OCR information into the LLM. For the visualization of qualitative examples, please refer to Appendix A.2.
Experiment Document-oriented VQA KIE
DocVQA VisualMRC Avg. DeepForm KLC Avg.
w/o CoT Pre-training 76.00 350.18 213.09 24.54 25.49 25.02
w/o CoT Annealing 77.40 353.08 215.24 34.23 26.21 30.22
w/o Patch Features 77.32 356.76 217.04 33.75 25.97 29.86
Random $VE$ $LE$ 77.42 355.54 216.48 31.31 25.76 28.54
DocLayLLM 77.43 358.68 218.06 35.41 26.25 30.83
Table 6: Ablation study on various design components of our DocLayLLM.
4.5 Ablation Study
We have also conducted additional experiments to explore the effectiveness of our proposed method. To conserve computational resources, we perform pre-training at just 25% of the standard scale. Our primary focus is to validate the structure of our design and the impact of the proposed CoT Pre-training and CoT Annealing techniques. The main results are presented in Table 6, where w/o CoT Pre-training and w/o CoT Annealing mean pre-training without CoT and supervised fine-tuning only with CoT data. w/o Patch features denotes the absence of patch features and the corresponding pre-training task, and Random $VE$ and $PE$ refers to the use of randomly initialized visual embedder and layout embedder.
The experimental results indicate that omitting CoT Pre-training leads to an obvious performance drop in both the Document-oriented VQA and KIE tasks, while not using CoT Annealing primarily affects the performance of the Document-oriented VQA task. Besides, excluding patch features or using randomly initialized $VE$ and $LE$ partially compromise the proposed extension, resulting in performance degradation in both the Document-oriented VQA and KIE tasks. Overall, the combined impact of these components ensures the efficiency and effectiveness of our method.
5 Conclusion
We propose DocLayLLM, an efficient and effective multi-modal extension of large language models for text-rich document understanding. By inserting document patch tokens and 2D positional tokens into the LLM through the simple linear transformation, it achieves an efficient multi-modal extension, enabling the LLM to perceive and process textual, visual, and layout information from the OCR results of the document. Additionally, we introduce CoT Pre-training and CoT Annealing techniques centered around CoT, further enhancing the LLM’s ability to handle OCR-derived information, thereby strengthening its understanding of text-rich documents. Despite using less data and fewer training resources, our method surpasses existing OCR-dependent TDU models, demonstrating strong zero-shot generalization capability. Besides, it also exceeds the performance of OCR-free TDU methods in the same multi-downstream-task fine-tuning setting, showcasing exceptional performance and further underscoring its efficiency and effectiveness.
For the future work, we aim to further unlock the power of DocLayLLM and explore involving more types of documents during our model learning, to develop a more versatile, robust, and powerful document understanding model.
Appendix A Appendix
A.1 More Implement Details
We implement our DocLayLLM using Llama2-7B-Chat (Touvron et al., 2023) and Llama3-8B-Instruct (Llama Team, 2024). The hyper-parameters for both pre-training and supervised fine-tuning are detailed in Table 7. As shown in the table, our DocLayLLM demonstrates efficiency, requiring fewer training resources while maintaining high performance. This underscores the method’s capability to deliver robust results without the need for extensive computational power, making it a resource-efficient solution for text-rich document understanding tasks.
Parameters Pre-Training Supervised Fine-Tuning
LoRA Rank 64 64
Batch Size 512 64
Max Length 2560 2560
Precision bf16 bf16
Trainable Parameters 170M/Llama2; 178M/Llama3 170M/Llama2; 178M/Llama3
Fixed Parameters 6.7B/Llama2; 8.0B/Llama3 6.7B/Llama2; 8.0B/Llama3
Learning Rate 1e-4 2e-5
Weight Decay 0.01 0.01
Scheduler cosine cosine
Adam Betas [0.9, 0.999] [0.9, 0.999]
Adam Epsilon 1e-8 1e-8
Epoch 1 3
Table 7: Hyper-parameters of DocLayLLM.
A.2 Qualitative Examples
We also provide some qualitative examples to visualize the performance of our DocLayLLM compared with other methods. Figure 2 shows the visualizations of our DocLayLLM, LayTextLLM (Lu et al., 2024), and DocOwl 1.5 (Hu et al., 2024a) on KIE benchmarks. Whether compared to OCR-free or OCR-dependent methods, it is evident that our DocLayLLM excels at handling documents with complex structures. Additionally, it more accurately locates answers and generates complete responses to questions.
We further present more visualizations across a wider range of benchmarks. In these comparisons, our DocLayLLM, using a portion of mixed downstream task data during the supervised fine-tuning process, is evaluated against the SOTA OCR-free method, which undergoes a similar fine-tuning process. As illustrated in Figure 3, the results highlight DocLayLLM’s superior capability in processing tables and its reduced hallucinations compared to DocOwl 1.5. This demonstrates the significant potential of our OCR-dependent approach in addressing text-rich document understanding tasks and underscores the rationale behind our proposed extension.
A.3 Preliminary Comparisons with MLLMs
Typically, many OCR-free TDU models are built upon an MLLM. With the rapid advancement of MLLMs, they have also shown impressive performance in text-rich document understanding. Therefore, we conduct a preliminary comparison between our DocLayLLM and some of the most powerful MLLMs available today, including the open-source InternVL2 (OpenGVLab, 2024) and the closed-source GPT-4o (OpenAI, 2024). These models are often evaluated on OCRBench (Liu et al., 2023b) to test their OCR capabilities. The OCRBench includes the KIE sub-task, consisting of 66 samples from FUNSD, 67 from SROIE, and 67 from POIE, amassing a total of 200 test samples for document question-answering. These are designed to evaluate the model’s key information extraction capabilities on text-rich documents. Thus, we compare our DocLayLLM against these models on the KIE sub-task in OCRBench.
Figure 2: Qualitative comparisons with OCR-free and OCR-dependent methods on KIE benchmarks, including FUNSD (Jaume et al., 2019), SROIE (Huang et al., 2019), and POIE (Kuang et al., 2023).
Figure 3: Qualitative comparisons with DocOwl 1.5 across various benchmarks. The document-oriented VQA tasks include DocVQA (Mathew et al., 2021), InfoVQA (Mathew et al., 2022), and VisualMRC (Tanaka et al., 2021); the KIE task includes DeepForm (Svetlichnaya, 2020); and the Table Understanding task includes WTQ (Pasupat and Liang, 2015). indicates that the model uses the training set corresponding to the evaluation set during training.
As shown in Table 8, even when compared with the most powerful MLLMs, our DocLayLLM, using significantly fewer training resources and a smaller parameter size, performs on par with or even outperforms them in TDU tasks. This highlights the robust performance of our proposed efficient and effective extension, further demonstrating the immense potential of OCR-dependent solutions in addressing text-rich document understanding challenges. It also underscores the significant promise of combining traditional OCR techniques with advanced LLMs.
Method KIE in OCRBench
FUNSD SROIE POIE In Total
InternVL2-8B (OpenGVLab, 2024) 52 49 59 160
GPT-4o (OpenAI, 2024) 53 55 61 169
DocLayLLMLlama3 (Ours) 62 58 58 178
Table 8: Comparisons with SOTA MLLMs on the KIE sub-tasks within OCRBench Liu et al. (2023b). The metric reflects the number of samples correctly answered.